1. (25 pts) Let F(x, y, z) = (2xy + 25)i + (4.r?y3 + 2yz?)j + (5.624 + 3y222)k and let C be the curve given parametrically byMay you please explain all steps? I want to understand this and am so confused.

Thanks!
Answer #1

1) we have F = (any + zs î+ (42} + 2y2²) Î + (5 x 2 + 3y 2) be given parametrically boy st) = (3t+1) î + t Î + stŔ for ostsi, (1,7) = 2423 side Integrate both we get g (1, 2) = (24 z3 dy h(2) 242 h(z) + 969,2) y² z² + h(z) i fry, z) = x²y 4 + x 2 +